what would 2/3, 50%, 62.5%, 0.6, and 6/8 from greatest to least.

Answer by LinnW(1048)

One way to work this is to represent each number as a decimal
with the same number of decimal positions. Let's use 3 decimal positions
So 2/3 50% 62.5% 0.6 6/8 become
0.667 0.500 0.625 0.600 0.750
So
2/3 = 0.667 approximately
50% = 0.500
62.5% = 0.625
0.6 = 0.600
6/8 = 0.750
So in order of lowest to highest we have
50% = 0.500
0.6 = 0.600
62.5% = 0.625
2/3 = 0.667
6/8 = 0.750
Please let me know if this is clear.